OpenStudy (anonymous):

Help? What is the simplified form of x over x squared minus y squared – y over x squared minus y squared ? 1 over x minus y 1 over x plus y negative 1 over x plus y 1 over y minus x

OpenStudy (kainui):

I suggest writing these all out and going from there.

OpenStudy (anonymous):

$$\frac{\frac{1}{x}-y^2-y}{x^2-y^2}$$

OpenStudy (anonymous):

Sorry..

OpenStudy (dumbcow):

ahhh no parenthesis ... what you wrote can be interpreted many ways can you use the equation editor or post a pic of the question?

OpenStudy (dumbcow):

\[\frac{x}{x^{2} -y^{2}}-\frac{y}{x^{2} -y^{2}} = \frac{x-y}{x^{2} -y^{2}}= \frac{x-y}{(x-y)(x+y)} = \frac{1}{x+y}\]